OpenStudy (xxlyh):

Two lines, C and D, are represented by the equations given below: Line C: y = x + 14 Line D: y = 3x + 2 Which of the following shows the solution to the system of equations and explains why? (6, 20), because both lines pass through this point (6, 20), because the point does not lie on any axis (3, 11), because one of the lines passes through this point (3, 11), because the point lies between the two axes

OpenStudy (anthonyym):

You can find the point from the given options. Plug them in to see which fits the equation. For example, for the point (6,20), x = 6 and y = 20

Mehek (mehek14):

Plug the value of y from the first equation for y in the second equation x + 14 = 3x + 2 Can you solve for x?

OpenStudy (xxlyh):

x = 6?

Mehek (mehek14):

Now plug 6 for x y = x + 14 y = 6 + 14 = ?

OpenStudy (xxlyh):

20

Mehek (mehek14):

So it's either A or B

OpenStudy (xxlyh):

I just wanna make sure because I'm taking a test xD

OpenStudy (xxlyh):

but thank you for helping me!

Mehek (mehek14):

Yw :)